The t statistic for a test of

?0:?=34

??:?>34
based on n = 17 observations has the value t = 1.32.

(a) What are the degrees of freedom for this statistic?

(b) Using the appropriate table in your formula packet, bound the p-value as closely as possible:
____ < p-value < ______
Answer #1

sal: Grisen Given that, He u = 3s Hai My 34 ( Right - tailed) t = 1:39 n=17 (a) The degree of freedom is - df=n-1 = 17-1 16,